The mean lifetime of 200 florescent light tubes gave a mean lifetime of 1560 hours with a standard deviation of 50hours, is it likely that a sample have come from a population with a mean lifetime of 1500 hours?
1
Expert's answer
2021-10-14T18:25:42-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1500"

"H_1:\mu\not=1500"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05, df=n-1"

"=200-1=199" ​​degrees of freedom, and the critical value for a two-tailed test is"t_c = 1.971957."

The rejection region for this two-tailed test is "R = \{t: |t| > 1.971957\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{1560-1500}{50/\sqrt{200}}=16.970563"

Since it is observed that "|t|= 16.970563>1.971957=t_c" it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value for two-tailed, "\alpha=0.05, df=199," "t=16.97" is "p\approx0," and since "p=0<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 1500, at the "\alpha = 0.05" significance level.